OpenStudy (anonymous):

A developer has land that has x feet of lake frontage. The land is to be subdivided into lots, each of which is to have either 80 feet or 100 feet of lake frontage. If 1/9 of the lots are to have 80 feet of frontage each and the remaining 40 lots are to have 100 feet of frontage each, waht is the value of x

OpenStudy (dumbcow):

40 lots is equal to 8/9 of N total lots 8/9N = 40 N = 40*9/8 = 45 so 40 have 100ft while 5 have 80ft x = 40*100 + 5*80 = 4400